Status Report

NASA Space Station On-Orbit Status 4 Feb 2004

By SpaceRef Editor
February 4, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 4 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below. 

The crew continued cargo transfers from the Progress resupply ship, assisted by a detailed Russian and US transfer/unpack list.   [As of last night, Progress was about 40% unloaded.  The crew has asked for weekly conferences to help manage the stowage of all the new items, which include ESA payloads for guest cosmonaut André Kuipers (the two payloads HEAT and ARGES are currently stowed in the Node).  A first 13P Unpack & Stowage conference with the ground is scheduled for tomorrow (2/5).  Progress M1-11 needs to be fully unloaded, restowed with trash and readied electronically for automated separation and departure prior to the upcoming Russian two-man EVA-9 on 2/26 (see below).]

CDR/SO Michael Foale tended the PromISS-3 (Protein Crystal Growth Monitoring by Digital Holographic Microscope 3) experiment in the MSG (Microgravity Science Glovebox), by removing videotape #2 and installing tape #3.   [The experiment involves a total of 12 tapes for recording the science.  Based on previous PromISS and PromISS-2 operations, the tape exchanges have been carefully planned to maximize tape utilization and minimize crew time.  The first six tapes are being exchanged approximately every 36 hours, the next five tapes will be exchanged every 72 hours, and the last tape will capture the remaining 6 days of science.  PromISS-3 software controls the recorder’s start and stop times (time lapse).]

Using the new procedure developed to prevent potential gas bubbles from getting into the Elektron-VM oxygen generator, Alexander Kaleri primed (filled) the KOV thermal loops’ EDV container with purified water from the multifiltration/purification column unit (BKO) while checking for any air bubbles.

On the Vozdukh carbon dioxide (CO2) removal system, Kaleri switched from the manual mode to automatic control for the air scrubber cycle between the two absorbent beds, in order to compensate for deactivation of the on-board timer.   [On the PU SOA (Air Scrubbing System Control Panel), Sasha set the vacuum pump cycle to 1 minute, CO2 max partial pressure to 3.5 mmHg, and airflow to 80%.].

The FE also updated the on-board procedures books with FDF (flight data file) enhancements and updates delivered by Progress 13P.   {About 350 pages were replaced with new sheets and three books added (two for EVA-9 tasks, one for 13P unloading/loading).]

Mike Foale conducted the daily routine maintenance of the SM’s SOZh life support system and the regular status checkup on autonomous Increment 8 payloads, while Sasha prepared the daily IMS (inventory management system) “delta” update file for automatic database export/import.

The crew worked out on TVIS treadmill, RED exerciser and VELO cycle with load trainer.

At 12:50am EST, the CDR set up and conducted an amateur radio pass with students at James F. Bay Elementary School, Seabrook, TX.   [The school is dedicated to providing excellent education with emphasis on mathematics and technology.]

Without requiring crew involvement, TsUP ran a test of the Progress M1-11’s sun sensor (SD) over Russian ground sites.

The station continues in solar-oriented XPOP attitude (x-axis perpendicular to orbit plane), with the solar array wings biased 45 degrees backward for drag reduction, until tomorrow (2/5), when it will maneuver to YVV (y-axis in velocity vector, i.e., flying “sidewise”) just after planned Russian thruster testing (at 6:02am EST).

Today at NASA JSC, the ISS Program Office is conducting the Soyuz 7S Delta Flight Readiness Review (FRR) on the special topic of the upcoming Russian Orlan-EVA-9.   [The EVA is scheduled for 2/26 (Thursday), with Foale and Kaleri egressing from the DC-1 docking module at ~4:00pm EST, for a duration of nominally 5h 32 min.  Considerable detail work has gone into preparations for the two-man spacewalk without intravehicular (IV) assistance, and some open work remains, including crew sleep shifting details, to be closed by 2/12.  An onboard training session of Orlan-suited ingress into the Soyuz 7S CRV (crew return vehicle) from the DC–1 (in the event of a leak of this module) is scheduled for 2/19, the second after the demo on 11/18/03.  Objectives of the spacewalk include installation of radiation measuring equipment (Matryoshka), photography of two Japanese contamination monitors (MPAC, SEED), removal and switching out of sampling panels (SKKs, removable cassette containers), replacement of a Russian contamination monitoring SKK (Kromka), ATV (automated transfer vehicle) support engineering (removal/relocation of LSV laser retroreflectors) for ESA, and removal of a foreign object (cable strap) from the WA-2 ham radio antenna.]

Also for EVA-9, on 2/10 the SSRMS (Space Station Remote Manipulator System) will be “walked off” onto the Lab PDGF (power & data grapple fixture) for covering the spacewalk with its video camera system.   [This activity will be preceded by procedural review by the crew and setup of the DOUG (dynamic operations ubiquitous graphics) with appropriate new software files.]

A detailed procedure is being developed for returning the space between the Lab science window’s two pressure panes (“Volume D”) to vacuum, in conjunction with the installation of the new flexhose jumper delivered on 13P.   [The depressurization will be done very slowly with a QD (quick-disconnect) throttling technique, designed to avoid flash condensation of any humidity in Volume D.  The crew will first practice the throttling technique by using male and female NOAPs (nitrogen/oxygen pressure adapters).]

Battery #8 in the SM has now been declared failed. An R&R (removal & replacement) will be scheduled soon for this 800A unit.   [This is the old SM battery #7, which was swapped with the original #8 to test ZRU-7 (charge/discharge unit #7).  ZRU-7 was found to be fully operational.]

In gratitude for the crew’s great work with the recent EarthKAM activity, enthusiastic middle school kids from Brunswick, MN, uplinked group photographs of themselves to Mike and Sasha.

Today’s CEO (crew earth observations) targets, in the current XPOP attitude constrained by flight rule to fewer near-vertical targets due to Lab window shutter closure and current condensation-prevention plan (limited to 90 min. in 24 hours), were Cape Town, South Africa (the city and its famous promontories lied just to the right of track as ISS approached the coast from the SW), Congo-Zimbabwe Biomass Burning (although this is not burning season in southern Africa, the crew was asked to take advantage of this break in the weather to locate and document burn scars from previous seasons, looking primarily left of track), Rio de Janeiro, Brazil (looking for Brazil’s largest and most famous city as it sprawls along the coast just left of track), and Tuamotu Archipelago (400mm-lens:  ISS pass was over the southeastern end of this extensive archipelago.  As the station approached from the SW, the crew was asked to try for near-nadir views with the long lenses to capture details of the coral reef structures of these islands).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:01pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On.  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off; SKV-2 is Off (repair incomplete).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.6; ppO2 (mmHg) — 159.5; ppCO2 (mmHg) — 3.9;
  • SM Transfer Compartment:  Pressure (mmHg/psi) — 761; temperature (deg C) — 18.6.
  • FGB Cabin:  Pressure (mmHg/psi) — 756; temperature (deg C) — 23.3.
  • Node:  Pressure (mmHg/psi) — 754.07; temperature (deg C) — 24.6 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg/psi) — 756.13; temperature (deg C) — 25.1; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg/psi) — 756.33; temperature (deg C) — 26.7; shell heater temp (deg C) — 24.4, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 25.1
  • PMA-2:  Shell heater temp (deg C) — 10.3.

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (suntracking) and bias-angled 45 deg. for drag reduction. 
  • SM batteries:  Battery #8, formerly known as #7, is failed (to be replaced); battery #4 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); battery #5 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 3482 kg (7676 lb) as of 1/29/04  [SM(755) + FGB(2727) + Progress M(0) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1/LEE B, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:45am EST [= epoch]):

  • Mean altitude — 368.0 km
  • Apogee — 373.2 km
  • Perigee — 362.8 km
  • Period — 91.9 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0007716
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29734

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.